Solve for x.
8400 = 2400(1.08^(x))
*its 1.08 to the x power*

Answer by LinnW(1048)

8400 = 2400(1.08^(x))
take log of each side
log(8400) =log(2400*(1.08^x))
log(8400) = log(2400) + log(1.08^x)
log(8400) = log(2400) + x*log(1.08)
add -log(2400) to each side
log(8400) - log(2400) = x*log(1.08)
divide each side by log(1.08)
(log(8400)-log(2400))/log(1.08) = x
x = approximately 16.2779
Check by entering 2400*(1.08^16.2779) into your calculator